OpenStudy (anonymous):

Let r(t) = − e^(6t) j + e^(–6t) k . Show that r(t) is parallel to r′′(t) , for any time t.

OpenStudy (anonymous):

r(t)=-e^(6t)j+e^(-6t)k=A r'(t)=-6e^(6t)j-6e^(-6t)k r''(t)=-36e^(6t)j+36e^(-6t)k=B A*B=36e^(36t^2)+36e^(36t^2)=72e^(36t^2) ||A||=sqrt(e^36t^2)+e^(36t^2))=sqrt(2e^(36t^2)) ||B||=sqrt(36^2e^(36t^2)+36^2e^(36t^2)) ||B||=sqrt(2592e^(36t^2)) ||A||||B||=72e^(36t^2) LET W=36t^2 A*B/||A||||B||=cos(x) 72e^w/72e^w=1 1=cos(x) cos^-1(1)=x x=2pi or 0 which concludes these two are parallel

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

welcom